OpenStudy (anonymous):

yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: Number of pages: 50|100|150|200 Cost : 1.25|1.85|2.45|3.05|Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.012x + 49.985 B. y = 0.012x + 1.25 C. y = 83.3x + 0.65 D. y = 0.012x + 0.65

OpenStudy (anonymous):

y = 0.012x + 0.65 ?

OpenStudy (zzr0ck3r):

first find the slope \(\frac{1.25-1.85}{50-100}=0.012\)

OpenStudy (zzr0ck3r):

now use any point to find the y intercept use point (50,1.25) and we get y = 0.012x+b becomes 1.25=0.012(50)+b so b = .65 and you are \(\huge CORRECT \color{green}{\checkmark}\)